An immersion heater supply heat at the rete of 500js-';

1,What is the quantity of heat produced in 25 minutes

2, If all the heat energy is given to a block of brass of mass 3kg initially at 30 degree celcius, what is the rise in temperature? (S.h.c.of brass is 380 jkg-'k-')


1
Expert's answer
2021-03-09T15:26:33-0500

1)

"Q=Pt=500\ \dfrac{J}{s}\cdot25\cdot60\ s=7.5\cdot10^5\ J."

2)

"Q=mc\Delta T,""\Delta T=\dfrac{Q}{mc}=\dfrac{7.5\cdot10^5\ J}{3\ kg\cdot380\ \dfrac{J}{kg\cdot\!^{\circ}C}}=658^{\circ}C.""\Delta T=T_f-T_i,""T_f=\Delta T+T_i=658^{\circ}C+30^{\circ}C=688^{\circ}C."